Home‎ > ‎Crew 142‎ > ‎

1113-CommandersReport

Commander Report

11/13/2014

Digby Tarvin

 

Today started at 7:00am as usual, with the kitchen crew whipping up some French toast using the remainder of our powdered egg and the bread that Carmel had backed the previous day. The choice was appropriate as Cyprien, the French member of the crew, was chef for the day. At Breakfast Christiane informed us that she had had a night time visitor who had been rummaging around in her bag and had scattered a few things around the room. Either our rodents are becoming more bold, or the hab is haunted.

 

The main topic of conversation at the breakfast table was the temperature, as last night was quite cool, and we had discovered that the habs main heat source was not functioning. This made getting up in the morning a little harder than usual, but everyone was still up on time.

 

Today was another day with with visitors from Earth. ABC4 Utah came to Mars to report on our operations. They were due shortly after noon, so the morning was spent working in the Greenhab on both biology experiments and the solar cooker, as well as on preparation for the afternoon EVA.

 

Our visitors arrived just after lunch, and stayed till about 14:30 to see the EVA depart the airlock at 14:25.

 

EVA08 was to return to our first checkpoint from yesterday, at 2S

0518078 4252397, and consisted

of Carmel, Cyprien, Dario and Vince, with Christiane serving as habcom. The objective was to complete the sampling of several interesting sites that were observed yesterday but which we did not have time to investigate.

 

As Christiane and myself were left at the hab, and are both mission engineers, we spent the time investigating some of the technical issues that remained in the hab.

By the time the EVA returned,

we had found the problem with the main hab heating system, and also repaired the audio system in the hab living area, and even produced some experimental simulated raspberry ice cream.

 

The first EVA member exited the airlock back into the hab at 17:09 while the other three exercised the three ATVs which have not yet exhibited problems by driving them from the hab to the road and back. The latter group re-entered the hab at 17:19, in good time to start on the evening reports - but without the overcoats this evening.